OpenStudy (anonymous):

how do you solve 10A^5 B^2 C^9/6A B^5 C?

OpenStudy (saifoo.khan):

\[\frac{10A^5 B^2 C^9}{6A B^5 C}\] \[\frac{5A^4C^8}{4B^4}\]

OpenStudy (mathteacher1729):

"solve" means "find the value of a given variable in terms of the others." So like, "solve 2x + 1 = 5" makes sense... x = 2. Your question is asking you to "simplify". :) It's an important difference.

OpenStudy (anonymous):

sorry i meant simplify

OpenStudy (mathteacher1729):

No prob. :)

OpenStudy (anonymous):

how would you simplify it?

OpenStudy (mathteacher1729):

Scroll down to the table: http://www.mathsisfun.com/algebra/exponent-laws.html Those are the exponent rules. Here are LOTS of examples: http://tutorial.math.lamar.edu/Classes/Alg/RealExponents.aspx :)